OpenStudy (anonymous):

What is the simplify sum of square root 3+square root 12

OpenStudy (anonymous):

\[\sqrt{3}+\sqrt{12}\]

OpenStudy (anonymous):

\[\sqrt3+ \sqrt{12} = \sqrt3 + 2\sqrt3=3\sqrt3\]

OpenStudy (anonymous):

Put the problem into common terms (if possible, like in this case) and combine those like terms.